# Prompt-prompted Mixture of Experts for Efficient LLM Generation

### v2

@@ 5 Experiments
We showcase the superb performance of GRIFFIN on numerous tasks and models (Section 5.1) while achieving lower latency (Section 5.2), along with a study on several of its properties like sampling experts, sequence length scaling, and random inputs (Section 5.3). All experiments are run on NVIDIA L40 GPUs.
@@ 5.1 Performance
Using various models, we evaluate on several generation and classification tasks. For generation, we evaluate on XSum [NCL18], CNN/DailyMail [${}^{+}$16], COQA [RCM19], and SCROLLS QASPER [${}^{+}$21, ${}^{+}$22]. For classification, we evaluate on HellaSwag [${}^{+}$19], PIQA [${}^{+}$20], COPA [RBG11], ARC-Easy/Challenge [${}^{+}$18], and BoolQ [${}^{+}$19]. With the exception of XSum and CNN/DailyMail, we use LM Evaluation Harness for our experiments [${}^{+}$23]. Aside from comparing with the original LLM, we also compare GRIFFIN with a static sequence-level MoE based on neuron magnitudes. Similar to neuron magnitude pruning, this baseline selects experts based on neuron magnitudes in $\bm{W}_{1}$ for the generation phase but uses the entire FF blocks for prompting like GRIFFIN. In the case of GLU variants, the neuron-wise norms of $\bm{W}_{1}$ and $\bm{W}_{g}$ are elementwise multiplied to produce the pruning metric. As we will see, this straightforward baseline achieves great classification results but falters for generation.
As our method is designed specifically for generation, we alter classification evaluations to simulate generation. In typical classification tasks, LLMs do not enter the generative phase since the final token output of the prompt phase indicates the class. Consequently, directly applying GRIFFIN for classification tasks trivially yields the exact performance of the original model. Therefore, we treat all tokens but the final input token as the prompt. Then, the model is forced to go into the generation phase for one step to produce the class.
We start with a look into the relationship between the sparsity levels and performance degradation. This translates to varying $k$ when we select the top-$k$ of our statistic $\bm{s}$. To compare the performance degradation across multiple tasks, we plot the ratio of the final performance metrics between GRIFFIN and the full model in Figure 4. We see most of the performance is preserved at 50% FF sparsity in Llama 2 7B, Gemma 7B, and Mistral 7B. Different tasks have different tipping points where performance sharply drops, which may be related to the difficulty of the task [${}^{+}$24].
Table 2: Generation tasks XSum (1-shot), CNN/DailyMail (1-shot), CoQA (0-shot), SCROLLS QASPER (0-shot) at 50% FF sparsity. Magnitude neuron pruning fails in almost every case while GRIFFIN effectively preserves performance.
@@
Fixing FF sparsity to be 50%, we evaluate on more tasks and models. Table 1 and Table 2 show the performance of GRIFFIN on classification and generation, respectively. We see that magnitude neuron pruning achieves reasonable results for classification but completely annihilates the original performance in most generation settings. In contrast, GRIFFIN achieves not only better performance than the baseline in most scenarios, but also preserves most of or matches the original performance on all tasks.
@@ 5.2 Efficiency
We now present efficiency metrics of GRIFFIN. We collect synthetic datasets with samples having identical lengths and average results across samples. Like many other MoE methods, GRIFFIN is ideal for single sample inputs, such as in the case of personal devices, so we use batch size 1 for these experiments. We plan to extend our method to larger batch sizes in future work. Using Hugging Face implementations of Llama 2 13B and Gemma 7B at FP16 precision, we measure the latency in different scenarios on an NVIDIA L40 GPU.
@@
2048+2048 0.3 78.5 67.7 / 65.0 67.4 / 65.0
Recalling that our magnitude selection baseline is essentially neuron pruning at generation, this has the best possible speed-up since there is no MoE overhead per sample. From Table 3, GRIFFIN matches the best case, producing up to a 1.16$\times$ and 1.25$\times$ improvement in latency for long generation at 50% FF sparsity in Gemma 7B and Llama 2 13B, respectively. This illustrates that our method is as fast as a static neuron pruned LLM during generation while being adaptive to preserve the accuracy of the full model. In offloading settings with large models, our method has the potential to further accelerate inference. For a prompt, GRIFFIN essentially performs structured pruning on the massive network, and if this pruned model can fit on a single device, it will avoid offloading for the entirety of generation.
@@ 5.3 Ablations and Analysis
Sampling-based Selection.
@@
Prompt vs. Generation Length.
We find that GRIFFIN can potentially be made more robust for long generation by lengthening the prompt. To see this, we use language modeling on the concatenated version of WikiText to simulate generation. For a length $S$ input into the FF block, we designate the first $P$ tokens as the prompt and the last $G$ tokens as the generated portion such that $P+G=S$. The prompt partition is used to calculate our statistic $\bm{s}$ and determine the experts. The prompt partition uses the full FF block while the generation partition only uses the selected experts. When comparing the original model with GRIFFIN, we only compute the perplexity of the outputs from the generation partition since the other outputs will be identical. Based on Figure 5, GRIFFIN gets closer to the full model outputs when the prompt length increases and generation length decreases, meaning the difficulty with long generation can be suppressed with longer prompts.
Sparsity in Random Sequences.
@@
6 Conclusion
In this work, we have shown a special form of sparsity in FF layers and a simple method to exploit it. Flocking is a curious phenomenon present in many LLMs where tokens within a sequence activate at similar intensities. This structure motivated the design of GRIFFIN, a learning-free MoE selection mechanism to remove FF neurons during inference at the sequence level which preserves the full model’s performance on a large collection of classification and generative tasks at 50% FF sparsity while achieving lower latency. Furthermore, its applicability extends beyond just ReLU-based LLMs, allowing MoE adaptation to be possible for many more models. With its straightforward algorithm and no-cost deployment, GRIFFIN expands the accessibility of numerous LLMs for generative inference.
Acknowledgements
Appendix A Gating Metric
Appendix B More Flocking Examples

### v3

@@ 5 Experiments
We showcase the superb performance of GRIFFIN on numerous tasks and models (Section 5.1) while achieving lower latency (Section 5.2), along with a study on several of its properties like sequence length scaling and batching (Section 5.3). All experiments are run on NVIDIA L40 GPUs.
@@ 5.1 Performance
Using various models, we evaluate on several generation and classification tasks. For generation, we evaluate on XSum [NCL18], CNN/DailyMail [NZG+16], COQA [RCM19], and SCROLLS QASPER [DLB+21, SSI+22]. For classification, we evaluate on HellaSwag [ZHB+19], PIQA [BZB+20], COPA [RBG11], ARC-Easy/Challenge [CCE+18], and BoolQ [CLC+19]. With the exception of XSum and CNN/DailyMail, we use LM Evaluation Harness for our experiments [GTA+23]. Aside from comparing with the original LLM, we also compare GRIFFIN with a static neuron pruning method based on neuron magnitudes. Similar to neuron magnitude pruning, this baseline selects expert neurons based on neuron magnitudes in $\bm{W}_{1}$ for the generation phase but uses the entire FF blocks for prompting like GRIFFIN. In the case of GLU variants, the neuron-wise norms of $\bm{W}_{1}$ and $\bm{W}_{g}$ are elementwise multiplied to produce the pruning metric. As we will see, this straightforward baseline achieves great classification results but falters for generation. Another baseline we consider also uses the full model for the prompt and Wanda [SLBK23] in FF blocks for generation, using the prompt activations to prune FF weights. We refer to this adaptive unstructured pruning baseline as Adaptive Wanda. Note that Adaptive Wanda is completely unstructured and does not reduce the FF activation dimension at all.
As our method is designed specifically for generation, we alter classification evaluations to simulate generation. In typical classification tasks, LLMs do not enter the generative phase since the final token output of the prompt phase indicates the class. Consequently, directly applying GRIFFIN for classification tasks trivially yields the exact performance of the original model. Therefore, we treat all tokens but the final input token as the prompt. Then, the model is forced to go into the generation phase for one step.
We start with a look into the relationship between the sparsity levels and performance degradation. This translates to varying $k$ when we select the top-$k$ of our statistic $\bm{s}$. To compare the performance degradation across multiple tasks, we plot the ratio of the final performance metrics between GRIFFIN and the full model in Figure 4. We see most of the performance is preserved at 50% FF sparsity in Llama 2 7B, Gemma 7B, and Mistral 7B. Different tasks have different tipping points where performance sharply drops, which may be related to the difficulty of the task [YJL+24].
@@
Table 2: Generation tasks XSum (1-shot), CNN/DailyMail (1-shot), CoQA (0-shot), SCROLLS QASPER (0-shot) at 50% FF sparsity. Magnitude neuron pruning fails in almost every case while GRIFFIN and Adaptive Wanda effectively preserve performance, though unlike GRIFFIN, Adaptive Wanda is completely unstructured.
@@
[…]
@@ 5.2 Efficiency
We now present efficiency metrics of GRIFFIN. We collect synthetic datasets with samples having identical lengths and average results across samples. Like many MoE methods, GRIFFIN is ideal for single sample inputs since it is adaptive, such as in the case of personal devices, so we use batch size 1 for these experiments. Using Hugging Face implementations of Llama 2 13B and Gemma 7B at FP16 precision, we measure the latency in different scenarios on an NVIDIA L40 GPU.
@@
2048+2048 0.3 87.1 67.7 / 65.0 67.4 / 65.0
Recalling that our magnitude selection baseline is essentially neuron pruning at generation, this has the best possible speed-up since there is no expert neuron selection overhead per sample. From Table 3, GRIFFIN matches the best case, producing up to a 1.29$\times$ and 1.25$\times$ improvement in latency for long generation at 50% FF sparsity in Gemma 7B and Llama 2 13B, respectively. This illustrates that our method is as fast as a static neuron pruned LLM during generation while being adaptive to preserve the accuracy of the full model. In offloading settings with large models, our method has the potential to further accelerate inference. For a prompt, GRIFFIN essentially performs structured pruning on the massive network, and if this pruned model can fit on a single device, it will avoid offloading for the entirety of generation.
@@ 5.3 Ablations and Analysis
Prompt vs. Generation Length.
We find that GRIFFIN can potentially be made more robust for long generation by lengthening the prompt. To see this, we use language modeling on the concatenated version of WikiText to simulate generation. For a length $S$ input into the FF block, we designate the first $P$ tokens as the prompt and the last $G$ tokens as the generated portion such that $P+G=S$. The prompt partition is used to calculate our statistic $\bm{s}$ and determine the expert neurons. The prompt partition uses the full FF block while the generation partition only uses the selected neurons. When comparing the original model with GRIFFIN, we only compute the perplexity of the outputs from the generation partition since the other outputs will be identical. Based on Figure 5, GRIFFIN gets closer to the full model outputs when the prompt length increases and generation length decreases, meaning the difficulty with long generation can be suppressed with longer prompts.
Sharing Selected FF Neurons.
We further verify that GRIFFIN is simultaneously robust to batching and more performative than fixed neuron pruning methods using similar pruning metrics to $\bm{s}$ in (6). As such, GRIFFIN maintains its reliability as we extend beyond single inputs.
For static pruning, we test two methods. One is to use the example shot in the prompt to compute $\bm{s}$. Then, for all samples with the same shot, we select the same FF neurons for generation. With the other method, we can extend (6) such that $\bar{\bm{s}}$ can be a pruning metric for multiple inputs (i.e. the same FF neurons will be pruned). Suppose that $\bm{s}_{i}$ is the metric for sample $i$ with prompt length $S_{i}$. Then, we use
$\bar{\bm{s}}=\sum_{i}\frac{\bm{s}_{i}}{\sqrt{S_{i}}}$ (7)
to select FF neurons. Finding a global $\bar{\bm{s}}$ across all prompts in a dataset is the second method. Therefore, for a particular dataset, both methods have fixed expert neurons across samples.
To extend GRIFFIN to larger batch sizes, we use $\bar{\bm{s}}$ from (7) where the samples in a batch are aggregated. Thus, adaptability is maintained. From Table 4, GRIFFIN (even with batch size >1) achieves better performance than the static methods, reinforcing the performance benefit of adaptability. Interestingly, even though performance decays as we increase the batch size, the decay is slow.
Model Full Shot Global GRIFFIN (1) GRIFFIN (4) GRIFFIN (16)
Llama 2 7B 27.15 21.11 23.28 24.75 24.26 23.75
Gemma 7B 26.86 19.76 22.48 25.86 25.01 24.37
Table 4: Rouge-1 scores of 1-shot XSum with varying ways to use the same selected neurons across multiple samples. From left to right, the scores are the result of the full model, FF neurons selected based on the prompt, FF neurons selected based on the entire dataset, and GRIFFIN in batch sizes of 1, 4, and 16. All pruning methods use the full model for the prompt and 50% of the FF neurons during generation.
6 Conclusion
In this work, we have shown a special form of sparsity in FF layers and a simple method to exploit it. Flocking is a curious phenomenon present in many LLMs where tokens within a sequence activate neurons at similar intensities. This structure motivated the design of GRIFFIN, a learning-free adaptive structured pruning mechanism to remove FF neurons during inference at the sequence level which preserves the full model’s performance on a large collection of classification and generative tasks at 50% FF sparsity while achieving lower latency. Furthermore, its applicability extends beyond just ReLU-based LLMs. With its straightforward algorithm and no-cost deployment, GRIFFIN expands the accessibility of numerous LLMs for generative inference.
Acknowledgements
Appendix A Pruning Metric
Appendix B Sampling-based Selection
@@
Appendix C Sparsity in Random Sequences
@@
Appendix D Generation Examples
We show example generated text in Figure 8. Adaptive Wanda and GRIFFIN produce similar summaries compared to the those produced by the full models and the target.
Figure 8: Generation examples by Llama 2 7B and Gemma 7B, given a prompt [BBC16] from XSum (1-shot).
Appendix E More Flocking Examples